OpenStudy (anonymous):

for the curve y=4x^3 -3x^5 find all the points at which tangent passes through the origin

OpenStudy (anonymous):

can any one ans this question

OpenStudy (anonymous):

i am not able to understand

OpenStudy (anonymous):

can u explain me???????

OpenStudy (anonymous):

at a tangent to a any curve the derevative at that point is zero so d/dx=0,then u take the x value and subtitute them at the equation y=4x^3-3x^5.

OpenStudy (anonymous):

but then everything becomes zero

OpenStudy (anonymous):

but there are more than 1 point , how to identify the other

OpenStudy (anonymous):

thnx

OpenStudy (zarkon):

if you want the tangent line to pass through the origin then the answers are \[\pm\frac{\sqrt{6}}{3},0\]

OpenStudy (zarkon):

that is \[x=\pm\frac{\sqrt{6}}{3},0\]